A body of mass 7.5kg is to be pulled up along a place which is included at 30° to the horizontal. If the efficiency of the place is 75%,what is the minimum force required to pull the body up the plane?
1
Expert's answer
2020-12-14T12:14:29-0500
"\eta F=mg \sin{30}\\(0.75) F=(7.5)(9.8) \sin{30}\\F=49\ N"